OpenStudy (thecuteone25):

@qwertty123 please help

OpenStudy (thecuteone25):

deposit of $3500. If his interest is compounded quarterly, how much is in Jeffrey’s account after 2 years? Round your answer to the nearest cent. interest compounded quarterly: A equals P left parenthesis 1 plus r over 4 right parenthesis to the power of 4 t end exponent

OpenStudy (qwertty123):

Hello! I am here!

OpenStudy (thecuteone25):

lol

OpenStudy (qwertty123):

just plug the values into your formula , then let your calculator take over A = 3500(1 + .08/4)^8 = 3500(1.02)^8 = ...

OpenStudy (thecuteone25):

ok thanks qwertty

OpenStudy (qwertty123):

You welcome! :D

OpenStudy (thecuteone25):

:)